Suppose license plates for cars are made up of 3 uppercase letters (A - Z)
and by 4 digits (0 - 9).How many license plates can there be using the system above if:there is at least 1 odd digit?

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3835)


26 letters in the English alphabet
26^3 = 17,576 different 3 letter combos where the letters can repeat.

Consider the case all 4 numeric digits are even.
There are 5 even single digits (0,2,4,6,8)
There are 5^4 = 625 ways to form a four digit number that has all even digits.
This is out of 10^4 = 10,000 ways to form a four digit sequence. Zero can be at the front.
10,000 - 625 = 9375 ways to have a sequence of 4 digits where at least one of them is odd.


Recap:
17,576 ways to form the letters portion
9375 ways to form the numbers portion

So,
17576*9375 = 164,775,000
represents the number of possible license plates.


Answer: 164,775,000

Answer by ikleyn(53751)

.

The problem is posed/worded incorrectly.

OBVIOUSLY,  it assumes that there is some ordering of letters and digits,
like  " the letters go first "  or something of this kind.

It assumes,  but the given problem says nothing about this ordering.


Without such ordering,  the number of all possible plates is much greater.


And without such given info about ordering,  the solution should be different.